# Optimal Transport for $e/\pi^{0}$ Particle Classification in LArTPC Neutrino Experiments

### v1

@@ VII.2.1 K-Nearest Neighbors (kNN)
To estimate the best hyperparameter values, we implement 5-fold cross-validation on the randomly mixed training dataset $N=N_{p,train}+N_{e,train}=500$. We first fix a hyperparameter value $k$. We split $N$ into five separate sets, $N^{i}$ for $i=1,...,5$. We select one of these to be our validation set and the remaining four to be our training e.g. $N_{val}=N^{0}=100$ and $n=\sum_{j\neq i}N^{j}=400$. We store the results of the kNN, then repeat the process until we have cycled through all possible choices of $N_{val}=N^{i}$. We then average the resulting 5 sets of stored kNN results. We repeat this process for a range of hyperparameter values $k$, and the best performing hyperparameter is selected. Initial scans of the hyperparameter $k$ shows best performance around $k=9$, which is why for all energy bins we scan over the hyperparameter space in the range $k\in[7,11]$ with an increment size of $1$. The final results are obtained by splitting the $N_{p,test}+N_{e,test}$ into 5 sets of size 100. The model with the best hyperparameter $k$ is retrained using the full $N_{p,train}+N_{e,train}$ set of events and is then evaluated on the 5 test sets; the average and standard deviation of the results are reported.
@@ VIII Results
Overall, the best performing method has an accuracy ranging from 0.828 to 0.876 for events above 100 MeV. Performance is shown to be relatively stable across all energy bins, with a slight drop at higher energies. The first energy bin at $0.05-0.1$ GeV is also shown to have lower relative performance. This is to be expected since this energy range is below the rest mass of a $\pi^{0}$, meaning the $\pi^{0}$ events in this energy bin are missing a significant portion of their energy due to effects such as inefficiencies in the 3D reconstruction and a fraction of the charge escaping the detector boundary. Overall, all methods perform well, however post-processing with both kNN and SVM boosts performance. SVM in particular makes the greatest improvement. In-line with previous works [38, 33], interpretable ML methods are shown to make good use of the features of OT distances and yield more competitive performance. Using 2D projections as input for OT (abbreviated as 2DOT) is shown to perform better than 3DOT when using min(OT) as the post-processing method, which makes sense as min(OT) may be particularly sensitive to cases such as those discussed in Fig. 6. However, kNN and SVM are able to overcome this difference. In particular, 3DOT+SVM has the best performance across all energy bins. As we will see in Sec. VIII.1, all three methods with both 2D and 3D OT outperform traditional reconstruction in $e/\pi^{0}$ classification.
@@ IX Conclusion
There are a variety of future directions motivated by this work. The success of OT methods applied to $e/\pi^{0}$ particle classification suggests applying OT to other reconstruction problems in LArTPC experiments such as $e/\gamma$ separation. The computational complexity of standard OT distance calculations warrants exploring strategies such linearized optimal transport (LOT) [38] for better computational efficiency. Finally, it would be worthwhile to explore the prospects for integrating OT into the reconstruction pipeline of neutrino LArTPC experiments comprising the SBN program and DUNE.
@@ Acknowledgements
We acknowledge the MicroBooNE Collaboration for making publicly available the data sets [64, 65] employed in this work. These data sets consist of simulated neutrino interactions from the Booster Neutrino Beamline overlaid on top of cosmic data collected with the MicroBooNE detector [2017 JINST 12 P02017]. We would like to thank Giuseppe Cerati for helping interpret the MicroBooNE public datasets and Pedro Machado for suggesting 2D projections. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Numbers DE-SC-0023730 and DE-SC-0011702. Parts of this work were performed at the Kavli Institute for Theoretical Physics, supported by the National Science Foundation under Grant No. NSF PHY-1748958. JNH was supported by the National Science Foundation under Grant No. NSF PHY-1748958 and by the Gordon and Betty Moore Foundation through Grant No. GBMF7392.
@@ Appendix A Gromov-Wasserstein Limited Run
Gromov-Wasserstein distances [89] were initially developed to calculate OT distances between distributions over different underlying metric spaces. For example, the OT distance between a 2D and a 3D distribution. To achieve this it relies only on relative distance information within each metric space. Concretely, let $\mu=\sum_{i}a_{i}~{}\delta_{x_{i}}$ and $\nu=\sum_{j}b_{j}~{}\delta_{y_{j}}$ be two discrete probability distributions over the ground spaces $X$ and $Y$, respectively. Let $D_{i,k}:=d_{X}(x_{i},x_{k})$ where $d_{X}(\cdot,\cdot)$ is the Euclidean distance between points $x_{i},x_{k}$ in the ground space $X$. Similarly, we have $D_{j,l}:=d_{Y}(y_{j},y_{l})$. Then the $p$-Gromov-Wasserstein distance is given by
$GW_{p}(\mu,\nu)=\left(\underset{\gamma\in\Pi(\mu,\nu)}{\rm min}\sum_{i,j,k,l}\left|D_{i,k}-D_{j,l}\right|^{p}\gamma_{i,j}\gamma_{k,l}\right)^{\frac{1}{p}},$ (2)

### v2

@@ VII.2.1 K-Nearest Neighbors (kNN)
To estimate the best hyperparameter values, we implement 5-fold cross-validation on the randomly mixed training dataset $N=N_{\pi^{0},train}+N_{e,train}=500$. We first fix a hyperparameter value $k$. We split $N$ into five separate sets, $N^{i}$ for i=1,,5i=1,...,5. We select one of these to be our validation set and the remaining four to be our training e.g. $N_{val}=N^{0}=100$ and $n=\sum_{j\neq i}N^{j}=400$. We store the results of the kNN, then repeat the process until we have cycled through all possible choices of $N_{val}=N^{i}$. We then average the resulting 5 sets of stored kNN results. We repeat this process for a range of hyperparameter values $k$, and the best performing hyperparameter is selected. Initial scans of the hyperparameter $k$ shows best performance around $k=9$, which is why for all energy bins we scan over the hyperparameter space in the range $k\in[7,11]$ with an increment size of $1$. Similar to what was done with min(OT), the final results are obtained by splitting the $N_{\pi^{0},test}+N_{e,test}$ into 5 sets of size 100. The model with the best hyperparameter $k$ is retrained using the full $N_{\pi^{0},train}+N_{e,train}$ set of events and is then evaluated on the 5 test sets; the average and standard deviation of the results are reported.
@@ VIII Results
Overall, the best performing method has an accuracy ranging from 0.828 to 0.876 for events above 100 MeV. Performance is shown to be relatively stable across all energy bins, with a slight drop at higher energies. The first energy bin at 0.050.10.05-0.1 GeV is also shown to have lower relative performance. This is to be expected since this energy range is below the rest mass of a $\pi^{0}$, meaning the $\pi^{0}$ events in this energy bin are missing a significant portion of their energy due to effects such as inefficiencies in the 3D reconstruction and a fraction of the charge escaping the detector boundary. Overall, all methods perform well, however post-processing with both kNN and SVM boosts performance. SVM in particular makes the greatest improvement. In-line with previous works [57, 56], interpretable ML methods are shown to make good use of the features of OT distances and yield more competitive performance. Using 2D projections as input for OT (abbreviated as 2DOT) is shown to perform better than 3DOT when using min(OT) as the post-processing method, which makes sense as min(OT) may be particularly sensitive to cases such as those discussed in Fig. 6. However, kNN and SVM are able to overcome this difference. In particular, 3DOT+SVM has the best performance across all energy bins. As we will see in Sec. VIII.1, all three methods with both 2D and 3D OT outperform traditional reconstruction in $e/\pi^{0}$ classification.
@@ IX Conclusion
There are a variety of future directions motivated by this work. The success of OT methods applied to $e/\pi^{0}$ particle classification suggests applying OT to other reconstruction problems in LArTPC experiments such as $e/\gamma$ separation. The computational complexity of standard OT distance calculations warrants exploring strategies such linearized optimal transport (LOT) [57] for better computational efficiency. Finally, it would be worthwhile to explore the prospects for integrating OT into the reconstruction pipeline of neutrino LArTPC experiments comprising the SBN program and DUNE using a variety of charge readout methods. While this work has shown promising performance using wire-based TPC images, comparable performance is expected for strip-based detectors and OT may be even more suited to pixel-based readout thanks to its intrinsic 3D readout.
@@ Acknowledgements
We acknowledge the MicroBooNE Collaboration for making publicly available the data sets [98, 97] employed in this work. These data sets consist of simulated neutrino interactions from the Booster Neutrino Beamline overlaid on top of cosmic data collected with the MicroBooNE detector [2017 JINST 12 P02017]. We would like to thank Giuseppe Cerati for helping interpret the MicroBooNE public datasets and Pedro Machado for suggesting 2D projections. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Numbers DE-SC-0023730 and DE-SC-0011702. Parts of this work were performed at the Kavli Institute for Theoretical Physics, supported by the National Science Foundation under Grant No. NSF PHY-1748958. JNH was supported by the National Science Foundation under Grant No. NSF PHY-1748958, the Gordon and Betty Moore Foundation through Grant No. GBMF7392, and by a UCSB Chancellor’s postdoctoral fellowship.
@@ Appendix A Gromov-Wasserstein Limited Run
Gromov-Wasserstein distances [Mémoli_2011] were initially developed to calculate OT distances between distributions over different underlying metric spaces. For example, the OT distance between a 2D and a 3D distribution. To achieve this it relies only on relative distance information within each metric space. Concretely, let $\mu=\sum_{i}a_{i}~\delta_{x_{i}}$ and $\nu=\sum_{j}b_{j}~\delta_{y_{j}}$ be two discrete probability distributions over the ground spaces $X$ and $Y$, respectively. Let $D_{i,k}:=d_{X}(x_{i},x_{k})$ where $d_{X}(\cdot,\cdot)$ is the Euclidean distance between points $x_{i},x_{k}$ in the ground space $X$. Similarly, we have $D_{j,l}:=d_{Y}(y_{j},y_{l})$. Then the $p$-Gromov-Wasserstein distance is given by
[…]